# Quantum Internet: Technologies, Protocols, and Research Challenges

### v1

@@ 2.1 Basic Concepts
From eq 18, a simultaneous bell state measurement at repeater station $R_{x}$ i.e., on qubits $R_{x_{1}}$ and $R_{x_{2}}$ would result in either of four Bell states i.e., $|\phi^{+}\rangle^{R_{x_{1}}R_{x_{2}}},|\phi^{-}\rangle^{R_{x_{1}}R_{x_{2}}},|\psi^{+}\rangle^{R_{x_{1}}R_{x_{2}}}$, and $|\psi^{-}\rangle^{R_{x_{1}}R_{x_{2}}}$. Meanwhile, the collective state of qubits of Alice and Bob i.e., $A$ and $B$ would be in the state: $\frac{1}{\sqrt{2}}\left(\alpha|00\rangle+\beta|11\rangle\right)^{AB},\frac{1}{\sqrt{2}}\left(\alpha|00\rangle-\beta|11\rangle\right)^{AB},\frac{1}{\sqrt{2}}\left(\alpha|01\rangle+\beta|10\rangle\right)^{AB}$, and $\frac{1}{\sqrt{2}}\left(\alpha|01\rangle-\beta|10\rangle\right)^{AB}$. The measurement results obtained at repeater station $R_{x}$ are communicated to Alice or Bob via a classical channel. Depending upon these measurement results a single-qubit gate i.e., $I$ for $|\phi^{+}\rangle$, $\sigma_{z}$ for $|\phi^{-}\rangle$, $\sigma_{x}$ for $|\psi^{+}\rangle$, and $\sigma_{x}\sigma_{z}$ for $|\psi^{-}\rangle$ which gets the collective state of Alice and Bob qubit transformed to an EPR pair $|\phi^{+}\rangle$. This concludes the entanglement-swapping procedure as summarized in table 3.
Table 3: Bell state analysis: Entanglement swapping
@@
Let the initial EPR pair i.e., $|\phi^{+}\rangle=\frac{1}{\sqrt{2}}\left(|00\rangle+|11\rangle\right)$ with qubits 1 and 2 be distributed by two neighbouring nodes via a noisy quantum channel, which degrades the quality of the original pure quantum states, hence diminishing their fidelity compared to the original Bell state and leading to a mixed, i.e., non-pure, state. The resulting mixed Bell state after exposure to this channel can be written using the following Werner’s state [15]:
@@
To purify this EPR pair we share another EPR pair with qubits 3 and 4 of the same bell state as before. If node A has qubits 1 and 3 while node B has qubits 2 and 4, then CNOT quantum gates with sources qubits 1 and 2 as sources and qubits 3 and 4 as targets are operated. Node A measures the qubit 3 and Node B measures the qubit 4 and they exchange their measured results through the classical channel (two-way signalling). If the measurement results match then the EPR pair with qubits 1 and 2 is kept or else discarded. Due to the measurement, the EPR pair is no longer entangled as soon as qubits 3 and 4 are measured. However, by using this sacrificial EPR pair the fidelity of retained EPR pair is given by [16]:
@@
Figure 4: Using additional resources (qubits) to enhance resilience against errors. a) Entanglement purification (modified from [18]); b) Quantum error correction (modified from [22])
@@
A logical qubit is an abstract representation of quantum information encoded within a set of physical qubits, making it resistant to errors through redundancy. In this setup, each qubit in a standard entanglement swapping protocol is replaced by an array of $n$ physical qubits utilizing $n$-qubit repetition code. In our case (Figure 4b), we have $n=3$ corresponds to the use of a three-qubit repetition code.
@@ 2.2 Components of Quantum Internet
Figure 5: Entanglement generation schemes. a) MeetInTheMiddle b) SenderRecever c) MidPointSource
A quantum repeater may be equipped with a quantum memory, i.e., a device that can store quantum states as qubits for a limited amount of time, and may also include a quantum processor, i.e., equipment for the execution of local operations through quantum gates. The requirement for quantum memory in a quantum repeater depends on the intended application and the entanglement distribution protocols employed by the quantum network. For instance, in prepare-and-measure quantum key distribution (QKD) networks, quantum memories are not necessary, whereas more advanced quantum applications, such as distributed quantum computing, rely on quantum memory to function effectively. From a protocol perspective, the need for quantum memory also varies. In networks where entanglement generation and swapping proceed sequentially or with minimal delay to directly entangle end nodes, quantum memory support may not be essential. However, if the protocol is designed such that, following initial entanglement generation, some links must wait for entanglement swapping to proceed, then quantum memory support becomes crucial. Quantum memory is also indispensable for protocols that involve purification or quantum error correction. Similarly, the inclusion of a quantum processor depends on the network protocols. If certain protocols are implemented that leverage a quantum processor to optimize the performance of the quantum network, then it becomes necessary. In the absence of such protocols, a quantum processor may not be required. An entanglement generator can be part of a quantum repeater if the entanglement generation scheme follows either the MeetInTheMiddle or SenderReceiver configuration [25]. In MeetInTheMiddle scheme, entanglement generation occurs at the outer nodes, with one qubit of the entangled pairs from each outer node sent to an intermediate station where entanglement swapping takes place. In contrast, in the SenderReceiver scheme, entanglement generation also occurs at the outer nodes as in the previous case. However, one qubit of entangled pair from one outer node is sent directly to the other outer node, eliminating the need for an intermediate station. For the MidpointSource scheme, a separate entanglement generation component positioned at the midpoint of a quantum network link would be required as shown in Figure 5.
@@
Quantum links: The quantum links are the quantum and classical channels that connect two neighbouring quantum repeaters or a quantum repeater and a quantum device. These links can be realised through fibre optic cables (represented as solid lines) or via free-space communication (represented as dotted lines), commonly used in ground-satellite networks, as shown in Figure 6. Because of current technology limitations, the generation of EPR pairs between neighbouring components is a probabilistic process, whose success rate depends on the specific technology used and the distance between the nodes, but in general is rather small ($\ll 0.5$). This aspect highlights the inherent variability and complexity in the structure of quantum networks.
Network Controller: A network controller is a logically centralised entity responsible for overseeing a specific segment of the quantum network within its jurisdiction. It manages the communication of Bell state measurement results from quantum repeaters to the end nodes, facilitating the completion of the entanglement swapping process. Additionally, the network controller handles routing decisions for incoming requests, ensuring efficient network operation. The network controller is commonly found in many proposed architectures of quantum networks, even if it may be an obstacle to scalability as the network grows, both in geographical size and number of nodes. Therefore, decentralised alternatives are also under study to cover such use cases, even if they are currently less explored due to the additional complexity required. Only time will tell which approach will dominate the future of the quantum internet.
@@ 3.1 Enhanced security or privacy
The applications in this category essentially hinges on providing uncompromising security or privacy compared to its classical counterparts. As introduced in Section 2, these applications use the fundamental principles of quantum mechanics i.e., the no-cloning theorem and quantum entanglement.
The advancements in quantum computing pose a significant threat to traditional public key cryptographic systems such as RSA [26], Diffie-Hellman [27], and ECC [28, 29]. To address this challenge, two primary solutions are proposed. The first involves utilising quantum cryptography, specifically quantum key distribution (QKD) [30, 31, 32], which is one of its most successful applications. QKD protocols allow two parties—such as quantum devices within the quantum internet—to share symmetric secret keys securely. These keys can then be used with encryption algorithms to securely transmit messages. The second solution is the adoption of post-quantum cryptography [33], which employs classical systems designed to be secure against quantum attacks.
@@
Following the theme of enhancing security and privacy, quantum technology offers various applications in distributed systems related to consensus and verification. A prime example is the quantum Byzantine agreement, developed by Ben-Or et al. [36], which ensures that a group of participants can reach consensus on a bit value, robust against faulty or malicious behaviors. Building on this foundation of trust, quantum technology extends its utility to secure voting and surveying systems. As outlined by Vaccaro et al. [37], these systems allow participants to cast votes or respond to surveys with guaranteed anonymity, enhancing the integrity of collective decision-making processes. Moreover, the field of quantum cryptography has made significant strides in enabling secure multiparty communications. Quantum Conference Key Agreement (CKA), as discussed in recent studies by Hahn et al. [38] and Murta et al. [39], facilitates the establishment of a shared secret key among multiple parties, crucial for coordinating actions across different nodes without compromising security. Another innovative application is certified deletion, introduced by Broadbent et al. [40], which allows for the verifiable destruction of encrypted information, ensuring that once a ciphertext is deleted, it remains irrecoverable, even if the decryption key is compromised. Further expanding the horizon, quantum leader election protocols, such as those explored by Ganz et al. [41], enable a group of distant, mutually distrustful entities to democratically elect a leader, ensuring fairness and transparency in critical decision-making scenarios. Additionally, the integration of quantum technologies into broader infrastructures is underway, with applications in the Internet of Things (IoT) being actively researched by Rahman et al. [42], and prospective uses in future 6G networks as envisioned by Rozenman et al. [43]. These advancements illustrate the expanding role of quantum technology in shaping modern technological landscapes, promising unprecedented levels of security and efficiency in digital communications and beyond.
@@ 3.2 Enhanced computing
Another application of great potential interest is quantum federated learning[47]: each client trains their model on their local dataset using a quantum computer, thereby keeping their data private. Instead of sharing private data, clients only exchange their model’s weights. The global model is then trained by aggregating these weights from all participating clients. While this is done today by exchanging only classical data, there are huge opportunities that can be unlocked by entanglement via the quantum internet since this would remove the need for the data to be decoded and re-encoded at each iteration. Finally, we mention that solutions have been proposed to enhance the accessibility and efficiency of quantum computing resources for diverse applications. Currently, they require only classical communications. For instance, Quantum-as-a-Service [48, 49] involves a Quantum API Gateway that recommends the most suitable quantum computer for running a specific quantum service in real-time. We can speculate that, as the quantum internet is deployed, such systems will evolve into more sophisticated versions building on the end-to-end distribution of entanglement among remote nodes.
@@ 3.3 Specialised applications
Additionally, there is the application of enhanced-baseline length for telescopes [52, 53]. Traditional optical interferometers are limited in their resolution due to restricted baseline lengths, noise, and signal loss during photon transmission between telescopes. The use of quantum networks can potentially overcome these limitations, allowing for interferometers with arbitrarily long baselines and thus dramatically improving their resolving power.
@@ 4 The Quantum Internet
We now delve into the main topic of this work. In Section 4.1 we introduce a possible roadmap of the quantum internet. Then we proceed in a bottom-up manner by illustrating the state-of-the-art physical connectivity technologies available (Section 4.2), the quantum repeater (Section 4.3), the main quantum network algorithms and protocols (Section 4.4), and finally we provide an overview of the stack models of the quantum internet that have been proposed so far, in Section 4.5.
@@ 4.1 Stages of quantum internet
As with any developing technology especially as sophisticated as the quantum internet, the implementation is only possible in stages. The different stages of the quantum internet have been categorised by the amount of incremental functionality available to the quantum devices or end nodes in [54]. The summary of the stages of quantum internet development is given in Table 4.
Stage 1 of the quantum internet initiates the quantum internet with basic point-to-point quantum key distribution (QKD) setups that depend on trusted intermediate nodes for key relay and security [55, 56]. It primarily features networks where end-to-end quantum communication is absent, relying on secure key generation between adjacent nodes using trusted nodes. To bolster security against potentially untrusted nodes, measurement-device-independent QKD protocols are employed, enhancing security without relying on the trustworthiness of measurement devices [57]. At this stage, the lack of end-to-end qubit entanglement prevents support for distributed quantum computing or quantum sensing. However, progress toward these goals begins with the ability of nodes to prepare and transmit a one-qubit state to any other node in the network. The transmission and measurement processes leverage post-selection, wherein only successful events—those where qubits are detected and measured correctly—are retained. Undetected or ”lost” qubits are disregarded.
@@
Stage 2 of the quantum internet enables end-to-end entanglement without the need for post-selection during transmission or measurement as in the case of Stage 1 above. However, due to the absence of quantum memory in the network, the entanglement must be used immediately after its creation. This stage allows for the successful distribution of end-to-end entanglement with a probability approaching 1.
@@
Stage 5 of the quantum internet represents the realization of a fully developed quantum network, enabling the complete range of applications in both computation and communication. At this stage, fault tolerance is achieved for all available quantum memory qubits, unlocking the full potential of the quantum internet for diverse and robust applications.

### v2

@@ 2.1 Basic Concepts
From eq 20, a simultaneous bell state measurement at repeater station $R_{x}$ that is, on qubits $R_{x_{1}}$ and $R_{x_{2}}$ would result in either of four Bell states that is, $|\phi^{+}\rangle^{R_{x_{1}}R_{x_{2}}},|\phi^{-}\rangle^{R_{x_{1}}R_{x_{2}}},|\psi^{+}\rangle^{R_{x_{1}}R_{x_{2}}}$, and $|\psi^{-}\rangle^{R_{x_{1}}R_{x_{2}}}$. Meanwhile, the collective state of qubits of Alice and Bob that is, $A$ and $B$ would be in the state: $\frac{1}{\sqrt{2}}\left(\alpha|00\rangle+\beta|11\rangle\right)^{AB},\frac{1}{\sqrt{2}}\left(\alpha|00\rangle-\beta|11\rangle\right)^{AB},\frac{1}{\sqrt{2}}\left(\alpha|01\rangle+\beta|10\rangle\right)^{AB}$, and $\frac{1}{\sqrt{2}}\left(\alpha|01\rangle-\beta|10\rangle\right)^{AB}$. The measurement results obtained at repeater station $R_{x}$ are communicated to Alice or Bob via a classical channel. Depending upon these measurement results a single-qubit gate that is, $I$ for $|\phi^{+}\rangle$, $\sigma_{z}$ for $|\phi^{-}\rangle$, $\sigma_{x}$ for $|\psi^{+}\rangle$, and $\sigma_{x}\sigma_{z}$ for $|\psi^{-}\rangle$ which gets the collective state of Alice and Bob qubit transformed to an EPR pair $|\phi^{+}\rangle$. This concludes the entanglement-swapping procedure as summarized in table 3.
Table 3: Bell state analysis for Entanglement swapping: Bell state measurement at quantum repeater’s ($R_{x}$) qubits $R_{x_{1}}$ and $R_{x_{2}}$ (column: $R_{x_{1}}R_{x_{2}}$ results) collapses (projects) the combined Alice’s qubit $A$ and Bob’s qubit $B$ to several different bell states (column: $AB$ result). Hence, a single-qubit gate(s) is applied at Bob’s qubit $B$ (column: applied single-qubit gate at Bob (B)) to arrive at correct end-to-end entanglement state between Alice’s qubit $A$ and Bob’s qubit $B$ (column: final state ($AB$)).
@@
Let the initial EPR pair that is, $|\phi^{+}\rangle=\frac{1}{\sqrt{2}}\left(|00\rangle+|11\rangle\right)$ with qubits 1 and 2 be distributed by two neighbouring nodes via a noisy quantum channel, which degrades the quality of the original pure quantum states, hence diminishing their fidelity compared to the original Bell state and leading to a mixed, that is, non-pure, state. The resulting mixed Bell state after exposure to this channel can be written using the following Werner’s state [16]:
@@
To purify this EPR pair, we share another EPR pair with qubits 3 and 4 of the same bell state as before. If node A has qubits 1 and 3 while node B has qubits 2 and 4, then CNOT quantum gates with sources qubits 1 and 2 as sources and qubits 3 and 4 as targets are operated. Node A measures qubit 3, and Node B measures qubit 4, and they exchange their measured results through the classical channel (two-way signalling). If the measurement results match, then the EPR pair with qubits 1 and 2 is kept, or else it is discarded. Due to the measurement, the EPR pair is no longer entangled as soon as qubits 3 and 4 are measured. However, by using this sacrificial EPR pair the fidelity of retained EPR pair is given by [17]:
@@
Figure 4: Using extra qubit resources to boost error resilience. a) Entanglement purification: Each EPR pair is paired with an additional, sacrificial EPR pair to enhance its fidelity (modified from [19]); b) Quantum error correction: Each EPR pair is supported by extra qubits arranged as encoded EPR pairs to reduce the error rate (modified from [25])
@@
A logical qubit is an abstract representation of quantum information encoded within a set of physical qubits, making it resistant to errors through redundancy. In this setup, each qubit in a standard entanglement swapping protocol is replaced by an array of $n$ physical qubits utilizing $n$-qubit repetition code. In our case (Figure 4b), we have $n=3$, which corresponds to the use of a three-qubit repetition code.
In the broader context of error correction, it is also instructive to introduce the general notation $[[n,k,d]]$, which succinctly characterizes an error-correcting code. Here, $n$ denotes the total number of physical qubits used in the encoding, $k$ represents the number of logical qubits encoded, and $d$ is the code distance—the minimum number of physical qubit errors required to cause an undetectable error. This distance determines the error-correcting capability of the code; specifically, a code with distance $d$ can detect up to $d-1$ errors and correct up to $\lfloor(d-1)/2\rfloor$ errors. This notation mirrors the classical code notation $[n,k,d]$. In our example utilizing the three-qubit repetition code, the parameters can be expressed as $[[3,1,3]]$; a single logical qubit is redundantly encoded into three physical qubits, and the code is capable of correcting any single-qubit error.
@@ 2.2 Components of Quantum Internet
Figure 5: Entanglement generation schemes. a) MeetInTheMiddle: Each of two distant stations holds one half of an EPR pair. The entangled qubits are sent to an intermediate station, where entanglement swapping establishes a direct end-to-end entanglement between the remote stations; b) SenderRecever: An entangled qubit generated at the source station is transmitted to the destination station. Here, entanglement swapping between this qubit and one from the destination’s EPR pair creates a direct entanglement link between the stations; c) MidPointSource: An intermediate station produces an EPR pair and sends one qubit to the source station and the other to the destination station. At each station, entanglement swapping between the received qubit and a qubit from the local EPR pair sets up end-to-end entanglement.
A quantum repeater may be equipped with a quantum memory, that is, a device that can store quantum states as qubits for a limited amount of time, and may also include a quantum processor, that is, equipment for the execution of local operations through quantum gates. The requirement for quantum memory in a quantum repeater depends on the intended application and the entanglement distribution protocols employed by the quantum network. For instance, in prepare-and-measure quantum key distribution (QKD) networks, quantum memories are not necessary, whereas more advanced quantum applications, such as distributed quantum computing, rely on quantum memory to function effectively. From a protocol perspective, the need for quantum memory also varies. In networks where entanglement generation and swapping proceed sequentially or with minimal delay to directly entangle end nodes, quantum memory support may not be essential. However, if the protocol is designed such that, following initial entanglement generation, some links must wait for entanglement swapping to proceed, then quantum memory support becomes crucial. Quantum memory is also indispensable for protocols that involve purification or quantum error correction. Similarly, the inclusion of a quantum processor depends on the network protocols. If certain protocols are implemented that leverage a quantum processor to optimize the performance of the quantum network, then it becomes necessary. In the absence of such protocols, a quantum processor may not be required. An entanglement generator can be part of a quantum repeater if the entanglement generation scheme follows either the MeetInTheMiddle or SenderReceiver configuration [28]. In the MeetInTheMiddle scheme, entanglement generation occurs at the outer nodes, with one qubit of the entangled pairs from each outer node sent to an intermediate station where entanglement swapping takes place. In contrast, in the SenderReceiver scheme, entanglement generation also occurs at the outer nodes, as in the previous case. However, one qubit of the entangled pair from one outer node is sent directly to the other outer node, eliminating the need for an intermediate station. For the MidpointSource scheme, a separate entanglement generation component positioned at the midpoint of a quantum network link would be required, as shown in Figure 5.
@@
Quantum links: The quantum links are the quantum and classical channels that connect two neighbouring quantum repeaters or a quantum repeater and a quantum device. These links can be realised through fibre optic cables (represented as solid lines) or via free-space communication (represented as dotted lines), commonly used in ground-satellite networks, as shown in Figure 6.
Because of current technology limitations, the generation of EPR pairs between neighbouring components is a probabilistic process, whose success rate depends on the specific technology used and the distance between the nodes, but in general is rather small ($\ll 0.5$).
This aspect highlights the inherent variability and complexity in the structure of quantum networks.
Network Controller: A network controller is a logically centralised entity responsible for overseeing a specific segment of the quantum network within its jurisdiction. It manages the communication of Bell state measurement results from quantum repeaters to the end nodes, facilitating the completion of the entanglement swapping process. Additionally, the network controller handles routing decisions for incoming requests, ensuring efficient network operation. The network controller is commonly found in many proposed architectures of quantum networks, even if it may be an obstacle to scalability as the network grows, both in geographical size and the number of nodes. Therefore, decentralised alternatives are also under study to cover such use cases, even if they are currently less explored due to the additional complexity required. Only time will tell which approach will dominate the future of the quantum internet.
@@ 3.1 Enhanced security or privacy
The applications in this category essentially hinge on providing uncompromising security or privacy compared to their classical counterparts. As introduced in Section 2, these applications use the fundamental principles of quantum mechanics, that is, the no-cloning theorem and quantum entanglement.
The advancements in quantum computing pose a significant threat to traditional public key cryptographic systems such as RSA [29], Diffie-Hellman [30], and ECC [31, 32]. To address this challenge, two primary solutions are proposed. The first involves utilising quantum cryptography, specifically quantum key distribution (QKD) [33, 34, 35], which is one of its most successful applications. QKD protocols allow two parties—such as quantum devices within the quantum internet—to share symmetric secret keys securely. These keys can then be used with encryption algorithms to transmit messages securely. The second solution is the adoption of post-quantum cryptography [36], which employs classical systems designed to be secure against quantum attacks.
@@
Following the theme of enhancing security and privacy, quantum technology offers various applications in distributed systems related to consensus and verification. A prime example is the quantum Byzantine agreement, developed by Ben-Or et al. [39], which ensures that a group of participants can reach consensus on a bit value, robust against faulty or malicious behaviors. Building on this foundation of trust, quantum technology extends its utility to secure voting and surveying systems. As outlined by Vaccaro et al. [40], these systems allow participants to cast votes or respond to surveys with guaranteed anonymity, enhancing the integrity of collective decision-making processes. Moreover, the field of quantum cryptography has made significant strides in enabling secure multiparty communications. Quantum Conference Key Agreement (CKA), as discussed in recent studies by Hahn et al. [41] and Murta et al. [42], facilitates the establishment of a shared secret key among multiple parties, crucial for coordinating actions across different nodes without compromising security. Another innovative application is certified deletion, introduced by Broadbent et al. [43], which allows for the verifiable destruction of encrypted information, ensuring that once a ciphertext is deleted, it remains irrecoverable, even if the decryption key is compromised. Further expanding the horizon, quantum leader election protocols, such as those explored by Ganz et al. [44], enable a group of distant, mutually distrustful entities to democratically elect a leader, ensuring fairness and transparency in critical decision-making scenarios. Additionally, the integration of quantum technologies into broader infrastructures is underway, with applications in the Internet of Things (IoT) being actively researched by Rahman et al. [45] and prospective uses in future 6G networks as envisioned by Rozenman et al. [46]. These advancements illustrate the expanding role of quantum technology in shaping modern technological landscapes, promising unprecedented levels of security and efficiency in digital communications and beyond.
@@ 3.2 Enhanced computing
Another application of great potential interest is quantum federated learning[50]: each client trains their model on their local dataset using a quantum computer, thereby keeping their data private. Instead of sharing private data, clients only exchange their model’s weights. The global model is then trained by aggregating these weights from all participating clients. While this is done today by exchanging only classical data, there are huge opportunities that can be unlocked by entanglement via the quantum internet since this would remove the need for the data to be decoded and re-encoded at each iteration. Finally, we mention that solutions have been proposed to enhance the accessibility and efficiency of quantum computing resources for diverse applications. Currently, they require only classical communications. For instance, Quantum-as-a-Service [51, 52] involves a Quantum API Gateway that recommends the most suitable quantum computer for running a specific quantum service in real time. We can speculate that, as the quantum internet is deployed, such systems will evolve into more sophisticated versions building on the end-to-end distribution of entanglement among remote nodes.
@@ 3.3 Specialised applications
Additionally, there is the application of enhanced-baseline length for telescopes [55, 56]. Traditional optical interferometers are limited in their resolution due to restricted baseline lengths, noise, and signal loss during photon transmission between telescopes. The use of quantum internet can potentially overcome these limitations, allowing for interferometers with arbitrarily long baselines and thus dramatically improving their resolving power.
@@ 4 The Quantum Internet
We now delve into the main topic of this work. In Section 4.1, we introduce a possible roadmap of the quantum internet. Then we proceed in a bottom-up manner by illustrating the state-of-the-art physical connectivity technologies available (Section 4.2), the quantum repeater (Section 4.3), the main quantum network algorithms and protocols (Section 4.4), and finally we provide an overview of the stack models of the quantum internet that have been proposed so far, in Section 4.5.
@@ 4.1 Stages of quantum internet
As with any developing technology, especially as sophisticated as the quantum internet, the implementation is only possible in stages. The different stages of the quantum internet have been categorised by the amount of incremental functionality
available to the quantum devices or end nodes in [57]. The summary of the stages of quantum internet development is given in Table 4.
Stage 1 of the quantum internet initiates the quantum internet with basic point-to-point quantum key distribution (QKD) setups that depend on trusted intermediate nodes for key relay and security [58, 59]. It primarily features networks where end-to-end quantum communication is absent, relying on secure key generation between adjacent nodes using trusted nodes. To bolster security against potentially untrusted nodes, measurement-device-independent QKD protocols are employed, enhancing security without relying on the trustworthiness of measurement devices [60]. At this stage, the lack of end-to-end qubit entanglement prevents support for distributed quantum computing or quantum sensing. However, progress toward these goals begins with the ability of nodes to prepare and transmit a one-qubit state to any other node in the network. The transmission and measurement processes leverage post-selection, wherein only successful events—those where qubits are detected and measured correctly—are retained. Undetected or “lost" qubits are disregarded.
@@
Stage 2 of the quantum internet enables end-to-end entanglement without the need for post-selection during transmission or measurement, as in the case of Stage 1 above. However, due to the absence of quantum memory in the network, the entanglement must be used immediately after its creation. This stage allows for the successful distribution of end-to-end entanglement with a probability approaching 1.
@@
Stage 5 of the quantum internet represents the realization of a fully developed quantum internet, enabling the complete range of applications in both computation and communication. At this stage, fault tolerance is achieved for all available quantum memory qubits, unlocking the full potential of the quantum internet for diverse and robust applications.